OpenStudy (anonymous):

Which statement best describes the effect on the graph of y = (x – 9)2 if the equation is changed to y = (x + 9)2? A.The graph moves up 18 units. B.The graph moves down 18 units. C.The graph moves left 18 units. D.The graph moves right 18 units.

OpenStudy (anonymous):

x-9=0 means x=9 x+9=0 means x=-9 what do you conclude?

OpenStudy (loser66):

he meant: if y = (x-9)^2, the value of y =0 is x =9 and the graph is